OpenStudy (anonymous):

If the sequence is geometric, what is its common ratio? 4, 8, 12, 16, ..

OpenStudy (anonymous):

is it +4 every time

OpenStudy (australopithecus):

an = 4n

OpenStudy (campbell_st):

the common ratio is found by comparing terms \[\frac{T_{2}}{T_{1}} = \frac{T_{3}}{T_{2}} = \frac{T_{4}}{T_{3}} ...\] if the sequence is geometric the cammon ratio should be 4 on every occassion

jimthompson5910 (jim_thompson5910):

yes it's +4 each time, so does that make it a geometric sequence?